OpenStudy (debpriya):

Find the remainder when 169 (144 to the power 25 ) is divided by 13 to the power 4. ( A. 5,021 B.3,927 C.4,563 D.3,251) ?

OpenStudy (stickman1998):

*hint* use a good calculator

OpenStudy (stickman1998):

Try Symbolab.com

OpenStudy (mathmath333):

is this the question \(\large \color{black}{\begin{align} & \dfrac{169\times144^{25}}{13^{4}} \hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (debpriya):

Yes yes

OpenStudy (mathmath333):

unfortunately there is no shortcut way i can find this can be calculated by only long way answer i think

OpenStudy (scooby6363):

144^25 = 9.1E53 13^4 = 28561 169 * 9.1E53 = 1.538E56 1.538E56 / 28561 = 5.385E51 or 5.385 * 10^51

OpenStudy (debpriya):

Could you please explain ?

jhonyy9 (jhonyy9):

like a first step rewrite the 169 = 13^2

jhonyy9 (jhonyy9):

so will get 13^2 *(12^2)^25 ---------------- so simplify by 13^2 and will get 13^4 12^50 ------ = ? 13^2

OpenStudy (debpriya):

Thank you !

OpenStudy (mathmath333):

but what after 12^50 ------ = ? 13^2